1. A vessel containing an unknown volume of argon gas in cm3 at 300°F and 101kPa was inverted and placed in cold ethanol. As the gas contracted, the pressure was quadrupled while the temperature reduced by one-third. What is the final volume of the gas in dm3.

2. In one sentence only, account for the differences in bond angle of the following isoelectronic species. CH4-109.5°, NH3-107°, H2O-105°
1
Expert's answer
2019-05-01T04:28:30-0400

1. a) pV = nRT

T1 = 300 0F = 422 K; T2 = 300/3 = 100 0F = 310.7 K.

p1V1 / T1 = p2V2 / T2

101000 x V1 / 422 = 404000 x V2 / 310.7

V2 = 0.001xV1

 

2. As H2O has two lone pairs so It repels the bond pairs much more and makes bond angle shorter of 104.5 degrees and as NH3 has one lone pair that repels the three bond pair but not much effectively and strongly as two lone pairs of water repel the bond pair so the bond angle between hydrogen atom of ammonia is 107.5 greater than that of water. Similarly an other molecule have sp3 hybridization that is methane CH4 molecule have no lone pair and each bond pair repels each other with equal force and bond angle between two adjacent hydrogen atoms becomes 109.5 degrees.